I'm not sure where this actually falls but my question is:
A certain kind of bacteria ha a growth constant of 0.04 (with respect to hours). If you start out with 120 grams, how much will you have after 7 hours?

Answer by ankor@dixie-net.com(22740)

A certain kind of bacteria ha a growth constant of 0.04 (with respect to hours). If you start out with 120 grams, how much will you have after 7 hours?
:
Using the formula:
A = Ao*(1.04)^t
Where
A = resulting amt
Ao = initial amt
t = time in hrs
;
A = 120(1.04)^7
A = 120 * 1.316
A = 157.9 grams